OpenStudy (anonymous):

Help y = -2x^2 - 12x + 1

Directrix (directrix):

What are the instructions?

OpenStudy (anonymous):

Identify the vertex for the graph of y = -2x^2 - 12x +1

Directrix (directrix):

Do you know how to begin in finding the vertex?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

im really bad at algebra so that is why i need help

OpenStudy (luigi0210):

I doubt you know derivatives? xD

OpenStudy (uri):

I was looking forward to solve a Quad equation......

OpenStudy (anonymous):

lol luigi i have no clue what that is

OpenStudy (mertsj):

The x coordinate of the vertex is: \[x=\frac{-b}{2a}\]

OpenStudy (mertsj):

After you find that, plug the answer into your equation and find y.

OpenStudy (anonymous):

You may want to get the answer by completing square. y = -2x^2 - 12x + 1 = -2 (x^2 + 6x ) + 1 = -2 [x^2 + 6x + (6/2)^2- (6/2)^2] + 1 = -2 (x^2 + 6x + 9 - 9) +1 = -2 (x^2 + 6x + 9 ) - 9(-2) + 1 = -2 (x^2 + 6x + 9 ) + 19 Then, reduce it into the form of y=a(x-h)^2+k The vertex is at (h,k)

OpenStudy (anonymous):

so the answer is -3,19